# thanks: Full author list given at the end of the paper

### v2

@@ I Introduction
In this paper, we observe for the first time the transverse polarization of the $\Lambda_{c}^{+}$ baryon in $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and obtain the phase difference ($\Delta\Phi$) between the EMFFs at thirteen CM energies ranging from $4600$ to $4951$ MeV, with an overall significance exceeding $10.0\sigma$. The decay asymmetry parameters ($\alpha_{BP}$ and $\Delta_{BP}$, which then leads to $\beta_{BP}$ and $\gamma_{BP}$) and phase shifts between the $S$- and $P$-wave ($\delta_{p}-\delta_{s}$) are measured for four CF channels: $\Lambda_{c}^{+}\to pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$. The decay of $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, notable for its clean structure and large branching fraction, is included to improve the transverse polarization measurement. By measuring the processes in samples separated by baryon charge, the CPC\!P asymmetry parameters, strong phase shifts and weak phase shifts between different partial waves are also extracted in the CF processes. This allows precise testing of the SM and provides an essential input for the theoretical predictions. The analysis is performed based on a multidimensional angular analysis of the cascade decays of $\Lambda_{c}^{+}$, as described in Ref. [15]. The charge conjugation channels are included in the analysis and are implied in the context unless explicitly stated otherwise.
@@ II Experimental apparatus and data sets
The Beijing Spectrometer III (BESIII) detector [48] records symmetric $e^{+}e^{-}$ collisions provided by the Beijing Electron Positron Collider (BEPCII) storage ring [49] in the CM energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [50, 51, 52]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps [53, 54, 55]. Data analyzed in this work were collected at thirteen energy points above the $\Lambda^{+}_{c}\bar{\Lambda}^{-}_{c}$ threshold. The integrated luminosity of the datasets analyzed in this work is estimated to be $6.4~\mathrm{fb}^{-1}$ [56, 57].
@@
Simulated data samples produced with a geant4-based [58] Monte Carlo package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [59, 60]. The inclusive simulation includes the production of open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc [59, 60]. All particle decays are modelled with evtgen [61, 62] using branching fractions either taken from the Particle Data Group (PDG) [41], when available, or otherwise estimated with lundcharm [63, 64]. Final state radiation (FSR) from charged final state particles is incorporated using the photos package [65]. The phase space (PHSP) signal simulation is generated by simulating all allowed kinematic configurations within the entire phase space for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, followed by the decay of $\Lambda_{c}^{+}$ into the signal final states and the inclusive decay of $\bar{\Lambda}_{c}^{-}$, which accounts for all possible decay modes. The signal simulation is based on the process model described in the helicity formalism with the decay asymmetry parameters measured in this work or cited from the PDG [41].
@@ III Measurement Method
III.1 Event selection and $\Lambda_{c}^{+}$ signal in the beam-constrained mass spectrum
A single tag approach [66] is employed, where only one $\Lambda_{c}^{+}$ or $\bar{\Lambda}_{c}^{-}$ baryon is reconstructed in each event. Four two-body hadronic decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$, and one three-body decay, $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, are studied in this research following Refs. [66, 67]. The intermediate states, $K_{S}^{0}$, $\Lambda$, $\Sigma^{0}$, and $\Sigma^{+}$ are reconstructed by combining $\pi^{+}\pi^{-}$, $p\pi^{-}$, $\gamma\Lambda$, and $p\pi^{0}$ candidates, respectively.
@@ III.1 Event selection and $\Lambda_{c}^{+}$ signal in the beam-constrained mass spectrum
Neutral showers are reconstructed in the EMC. Showers not associated with any charged track are identified as photon candidates. The deposited energy of each shower in the EMC must satisfy the following conditions: $E_{\text{barrel}}>25$ MeV for $|\cos\theta|<0.80$, $E_{\text{end-cap}}>50$ MeV for $0.86<|\cos\theta|<0.92$. The EMC time difference from the event start time is required to be less than 700 ns to suppress electronic noise and showers unrelated to the event. The $\pi^{0}$ candidates are reconstructed from photon pairs with invariant masses satisfying: $115~\text{MeV}/c^{2}<M(\gamma\gamma)<150~\text{MeV}/c^{2}$.
@@
To suppress the peaking backgrounds for some decay channels, like $\Lambda_{c}^{+}\to\Lambda l^{+}\nu_{l}(l=e/\mu)$ background in the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, and $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, additional requirements are applied. Positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. For the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, it is required that $\mathcal{L}^{\prime}(e)<0.05$ and $\mathcal{L}(\pi)>0.001$ are satisfied to suppress the $\Lambda_{c}^{+}\to\Lambda e^{+}\nu_{e}$ background. For the $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, the information from the muon counter detector is utilized, and the penetrated depth of the $\pi^{+}$ candidates is required to be less than $5~\text{cm}$. A kinematic variable, the energy difference $\Delta E\equiv E_{\Lambda_{c}^{+}}-E_{\text{beam}}$, is used to suppress background, where $E_{\Lambda_{c}^{+}}$ is the total measured energy of the $\Lambda_{c}^{+}$ candidate, and $E_{\text{beam}}$ is the beam energy. For the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, $pK^{0}_{S}$, $\Lambda\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays the requirements are $\Delta E\in(-29,26)$, $\Delta E\in(-21,18)$, $\Delta E\in(-23,21)$, and $\Delta E\in(-67,32)$, respectively. For the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, an additional photon can be selected from the process $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and its $\Delta E$ is expected to be larger than zero. The requirement $\Delta E\in(-33,20)~\text{MeV}$ is imposed and can suppress this type of background efficiently.
@@
A large background for the channel $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$ is from $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ combined with a random photon. This background, which may pose hidden dangers to the angular analysis, necessitates the application of a four-constraint kinematic fit to the $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$, where $\Sigma^{+}$ decays to proton and $\pi^{0}$, $\pi^{0}\to\gamma\gamma$. The quality of the fit, $\chi^{2}$, is calculated by constraining two $\pi^{0}$s, $\Sigma^{+}$, and $\Lambda_{c}^{+}$ to their world average masses quoted from PDG [41]. After optimization by $N^{2}_{S}/(N_{S}+N_{B})^{3/2}$ [68, 69], the requirement is set that the $\chi^{2}$ must be less than 12, where $N_{S}$ and $N_{B}$ represent the signal and background yields, respectively.
@@
After applying the above selection criteria, the beam-constrained mass $M_{\text{BC}}\equiv\sqrt{E_{\text{beam}}^{2}/c^{4}-|\vec{p}_{\Lambda_{c}^{+}}|^{2}/c^{2}}$, is defined to identify signal candidates, and $\vec{p}_{\Lambda_{c}^{+}}$ is the momentum of the $\Lambda_{c}^{+}$ candidates calculated in the $e^{+}e^{-}$ rest frame. An unbinned maximum-likelihood fit is performed on the distribution of $M_{\text{BC}}$. The shapes of both correctly and mis-reconstructed signal events are modeled with the simulation, convolved with a Gaussian function to account for the resolution difference between the data and the simulation. The truth matching technique classifies events as correctly reconstructed if the angle between the reconstructed and true photon momentum directions from the simulation is below 10 degrees; otherwise, they are mis-reconstructed. The background shapes are described by an ARGUS function [70], except in the case of the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel at CM energies larger than $4740$ MeV, where a second-order Chebyshev polynomial function is employed to model the background. The fit results are shown in Fig. 2 for all channels at $4600$ MeV. The obtained signal yields, combining all energy points, are $50083\pm 233$, $9619\pm 96$, $5742\pm 74$, $2487\pm 48$ and $1268\pm 37$ in the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$$pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$ channels, respectively, within the signal region of $M_{\text{BC}}\in[2.282,2.291]~\text{GeV}/c^{2}$.
@@
Figure 2: Beam-constrained mass spectrum for $\Lambda_{c}^{+}$ signal at $4600$ MeV. Each panel shows a different signal channel as labeled. The Black points with error bars are data. The red, and black shaded region indicate the fitted signal and mis-reconstructed signal events. The yellow region represents the fitted combinatorial background.
@@ III.2 Helicity system and angular distribution definitions
A helicity system and helicity angles are employed in constructing the amplitude for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, $\Lambda_{c}^{+}\to BP$. Taking the channel $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ as an example, a kinematic diagram of this decay is shown in Fig. 3. For the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, the momenta of the $\Lambda_{c}^{+}$ and $\bar{\Lambda}_{c}^{-}$ baryons are defined in the CM system of the $e^{+}e^{-}$ pair, and $\theta_{0}$ is the polar angle of the $\Lambda_{c}^{+}$ in CM system. For the helicity system describing $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ decay, the momenta of the $\Lambda$ and the $\pi^{+}$ are defined in the $\Lambda_{c}^{+}$ frame. The angle $\phi_{1}$ is that between the $e^{+}\Lambda_{c}^{+}$ and $\Lambda\pi^{+}$ planes, and $\theta_{1}$ is the polar angle of the $\Lambda$ momentum defined in the rest frame of the $\Lambda_{c}^{+}$ baryon. For the helicity system describing the $\Lambda\to p\pi^{-}$, $\phi_{2}$ and $\theta_{2}$ adopt similar definitions. Then the formula of angular distributions related to this helicity frame based on the language of helicity amplitudes and Wigner-D function can be derived [15].
@@
The joint angular distribution of the single-side decay chains can be derived from polarization transmission [3] or equivalently the helicity density matrix [71]. The angular distribution of the decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to\Lambda(\to p\pi^{-})\pi^{+})$ is shown below:
@@
where the $\alpha_{p\pi^{-}}$ denotes the decay asymmetry parameter in the weak hadronic decay $\Lambda\to p\pi^{-}$. The parameter $\Delta_{\Lambda\pi^{+}}$ is the phase shift between the helicity amplitudes of $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ satisfying $\beta_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\sin\Delta_{\Lambda\pi^{+}},\gamma_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\cos\Delta_{\Lambda\pi^{+}}$. Formulas of other two-body decays can be found in Ref. [15].
Although the analysis of the angular distribution of these four two-body CF channels provides effective constraints, the three-body channel $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, with a low background and large production rate, brings beneficial additional sensitivity. Benefiting from the transition amplitude analysis reported by the LHCb Collaboration [72], the angular distribution of decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to pK^{-}\pi^{+})$ can be expressed explicitly as:
@@
where $\phi_{1},\theta_{1},\phi_{2}$ are the helicity angles for the intermediate state ($K^{-}\pi^{+}$), $R_{ij}(\phi_{1},\theta_{1},\phi_{2})$ represents the 3-dimension $Z$-$Y$-$Z$ rotation matrix and $\kappa$ denotes a set of Dalitz-plot variables $(M(pK^{-})^{2},M(K^{-}\pi^{+})^{2})$. The $\vec{\alpha}(\kappa)$ and $I_{0}(\kappa)$ are model-agnostic representations for polarization dependence of the decay rate and the total differential decay rate over $\kappa$, respectively, which are taken from the $\Lambda_{c}^{+}$ polarimetry with the amplitude model both measured by LHCb Collaboration [72, 73].

### v3

@@ I Introduction
In this paper, we observe for the first time the transverse polarization of the $\Lambda_{c}^{+}$ baryon in $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and obtain the phase difference ($\Delta\Phi$) between the EMFFs at thirteen CM energies ranging from $4600$ to $4951$ MeV, with an overall significance exceeding 10.0 standard deviations. The decay asymmetry parameters ($\alpha_{BP}$ and $\Delta_{BP}$, which then leads to $\beta_{BP}$ and $\gamma_{BP}$) and phase shifts between the $S$- and $P$-wave ($\delta_{p}-\delta_{s}$) are measured for four CF channels: $\Lambda_{c}^{+}\to pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$. The decay of $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, notable for its clean structure and large branching fraction, is included to improve the transverse polarization measurement. Unless otherwise specified, the charge-conjugate channels are included throughout this article. By measuring the processes in samples separated by baryon charge, the CPC\!P parameters, strong phase shifts and weak phase shifts between different partial waves are also extracted in the CF processes. This allows precise testing of the SM and provides an essential input for the theoretical calculations. The phase differences $\delta_{p}-\delta_{s}$ and $\Delta\Phi$ encode the characteristics of the $\Lambda_{c}^{+}$ decay and production processes, respectively. Besides the involved strong interaction, the former is associated with the weak interaction, while the latter connects the electromagnetic one. CPC\!P violation can arise only from the weak interaction accompanying with the strong interaction. The analysis is performed based on a multidimensional angular analysis of the cascade decays of $\Lambda_{c}^{+}$, as described in Ref. [BESIII:2019odb].
@@ II Experimental apparatus and data sets
The Beijing Spectrometer III (BESIII) detector [BESIII:2009fln] records symmetric $e^{+}e^{-}$ collisions provided by the Beijing Electron Positron Collider (BEPCII) storage ring [Yu:2016cof] in the CM energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [BESIII:2020nme, Lu2020-hf, Zhang2022-id]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps [Guo2017-of, Li2017-zt, Cao2020-vl]. Data analyzed in this work were collected at thirteen energy points above the $\Lambda^{+}_{c}\bar{\Lambda}^{-}_{c}$ threshold. The integrated luminosity of the datasets analyzed in this work is estimated to be $6.4~\mathrm{fb}^{-1}$ [BESIII:2022dxl, BESIII:2022ulv].
@@
Simulated data samples produced with a geant4-based [GEANT4:2002zbu] Monte Carlo package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [Jadach:2000ir, Jadach:1999vf]. The inclusive simulation includes the production of open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc [Jadach:2000ir, Jadach:1999vf]. All particle decays are modelled with evtgen [Lange:2001uf, Ping:2008zz] using branching fractions either taken from the Particle Data Group (PDG) [ParticleDataGroup:2024cfk], when available, or otherwise estimated with lundcharm [Chen:2000tv, Yang:2014vra]. Final state radiation (FSR) from charged final state particles is incorporated using the photos package [Barberio:1990ms]. The phase space (PHSP) signal simulation is generated according to a pure phase-space model, assuming a uniform distribution of the available kinematic phase space without any dynamical amplitudes or polarization effects for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, followed by the decay of $\Lambda_{c}^{+}$ into the signal final states and the inclusive decay of $\bar{\Lambda}_{c}^{-}$, which accounts for all possible decay modes. The signal simulation is based on the process model described in the helicity formalism with the decay asymmetry parameters measured in this work or cited from the other measurements [BESIII:2022qax, ParticleDataGroup:2024cfk].
@@ III Measurement Method
III.1 Event selection and $\Lambda_{c}^{+}$ signal
in the beam-constrained mass spectrum
A single tag approach [BESIII:2017kqg] is employed, in which an event is accepted as an exclusive $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ final state when either a $\Lambda_{c}^{+}$ or its antiparticle is reconstructed in one of the decay channels studied, using all particles in the same side as the reconstructed baryon. Four two-body hadronic decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$, and one three-body decay, $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, are studied in this research following Refs. [BESIII:2017kqg, BESIII:2022xne]. The intermediate states, $K_{S}^{0}$, $\Lambda$, $\Sigma^{0}$, and $\Sigma^{+}$ are reconstructed by combining $\pi^{+}\pi^{-}$, $p\pi^{-}$, $\gamma\Lambda$, and $p\pi^{0}$ candidates, respectively.
@@ III.1 Event selection and $\Lambda_{c}^{+}$ signal
Neutral showers are reconstructed in the EMC. Showers not associated with any charged track are identified as photon candidates. The deposited energy of each shower in the EMC must satisfy the following conditions: $E_{\text{barrel}}>25$ MeV for $|\cos\theta|<0.80$, $E_{\text{end-cap}}>50$ MeV for $0.86<|\cos\theta|<0.92$. The EMC time difference from the event start time is required to be less than 700 ns to suppress electronic noise and showers unrelated to the event, where the event start time corresponds to the time at which the $e^{+}e^{-}$ collision occurs. The $\pi^{0}$ candidates are reconstructed from photon pairs with invariant masses satisfying: $115~\text{MeV}/c^{2}<M(\gamma\gamma)<150~\text{MeV}/c^{2}$.
@@
To suppress the peaking backgrounds for some decay channels, like $\Lambda_{c}^{+}\to\Lambda l^{+}\nu_{l}(l=e/\mu)$ background in the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, and $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, additional requirements are applied. Positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. For the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, it is required that $\mathcal{L}^{\prime}(e)<0.05$ and $\mathcal{L}(\pi)>0.001$ are satisfied to suppress the $\Lambda_{c}^{+}\to\Lambda e^{+}\nu_{e}$ background, which rejects roughly 95% of these backgrounds with only a negligible signal loss. For the $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, the information from the muon counter detector is utilized, and the penetrated depth of the $\pi^{+}$ candidates is required to be less than $5~\text{cm}$. These selections effectively reduce peaking backgrounds $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ by 80% while keeping over 80% of the signal events. A kinematic variable, the energy difference $\Delta E\equiv E_{\Lambda_{c}^{+}}-E_{\text{beam}}$, is used to suppress background, where $E_{\Lambda_{c}^{+}}$ is the total measured energy of the $\Lambda_{c}^{+}$ candidate, and $E_{\text{beam}}$ is the beam energy. For the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, $pK^{0}_{S}$, $\Lambda\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays the requirements are $\Delta E\in(-29,26)~\text{MeV}$, $\Delta E\in(-21,18)~\text{MeV}$, $\Delta E\in(-23,21)~\text{MeV}$, and $\Delta E\in(-67,32)~\text{MeV}$, respectively. For the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, an additional photon can be selected from the process $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and its $\Delta E$ is expected to be larger than zero. The requirement $\Delta E\in(-33,20)~\text{MeV}$ removes about 54% of this type of backgrounds while retaining more than 98% of signal events.
@@
A large background for the channel $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$ is from $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ combined with a random photon. This background, which may pose hidden dangers to the angular analysis, necessitates the application of a four-constraint kinematic fit to the $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$, where $\Sigma^{+}$ decays to proton and $\pi^{0}$, $\pi^{0}\to\gamma\gamma$. The quality of the fit, $\chi^{2}$, is calculated by constraining two $\pi^{0}$s, $\Sigma^{+}$, and $\Lambda_{c}^{+}$ to their world average masses quoted from PDG [ParticleDataGroup:2024cfk]. After optimization by $N^{2}_{S}/(N_{S}+N_{B})^{3/2}$ [LHCb:2021chn, LHCb:2022ogu], the requirement is set that the $\chi^{2}$ must be less than 12, where $N_{S}$ and $N_{B}$ represent the signal and background yields, respectively.
@@
After applying the above selection criteria, the beam-constrained mass $M_{\text{BC}}\equiv\sqrt{E_{\text{beam}}^{2}/c^{4}-|\vec{p}_{\Lambda_{c}^{+}}|^{2}/c^{2}}$, is defined to identify signal candidates, and $\vec{p}_{\Lambda_{c}^{+}}$ is the momentum of the $\Lambda_{c}^{+}$ candidates calculated in the $e^{+}e^{-}$ rest frame. An unbinned maximum-likelihood fit is performed on the distribution of $M_{\text{BC}}$. The shapes of both correctly and mis-reconstructed signal events are modeled with the simulation, convolved with a Gaussian function to account for the resolution difference between the data and the simulation. The truth matching technique classifies events as correctly reconstructed if the angle between the reconstructed and true photon momentum directions from the simulation is below 10 degrees; otherwise, they are mis-reconstructed. The background shapes are described by an ARGUS function [ARGUS:1990hfq], except in the case of the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel at CM energies larger than $4740$ MeV, where a second-order Chebyshev polynomial function is employed to model the background. The fit results are shown in Fig. 3 for all channels at $4600$ MeV. The obtained signal yields, combining all energy points, are $50083\pm 233$, $9619\pm 96$, $5742\pm 74$, $2487\pm 48$ and $1268\pm 37$ in the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$$pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$ channels, respectively, within the signal region of $M_{\text{BC}}\in[2.282,2.291]~\text{GeV}/c^{2}$.
@@
Figure 3: Beam-constrained mass spectrum for $\Lambda_{c}^{+}$ signal at $4600$ MeV. Each panel shows a different signal channel as labeled. The Black points with error bars are data. The black shaded region indicates the mis-reconstructed signal events. The green region represents the fitted combinatorial background.
@@ III.2 Helicity system and angular distribution definitions
A helicity system and helicity angles are employed in constructing the amplitude for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, $\Lambda_{c}^{+}\to BP$. Taking the channel $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ as an example, a kinematic diagram of this decay is shown in Fig. 4. For the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, the momenta of the $\Lambda_{c}^{+}$ and $\bar{\Lambda}_{c}^{-}$ baryons are defined in the CM system of the $e^{+}e^{-}$ pair, and $\theta_{0}$ is the polar angle of the $\Lambda_{c}^{+}$ in CM system. For the helicity system describing $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ decay, the momenta of the $\Lambda$ and the $\pi^{+}$ are defined in the $\Lambda_{c}^{+}$ frame. The angle $\phi_{1}$ is that between the $e^{+}\Lambda_{c}^{+}$ and $\Lambda\pi^{+}$ planes, and $\theta_{1}$ is the polar angle of the $\Lambda$ momentum defined in the rest frame of the $\Lambda_{c}^{+}$ baryon. For the helicity system describing the $\Lambda\to p\pi^{-}$, $\phi_{2}$ and $\theta_{2}$ adopt similar definitions. Then the formula of angular distributions related to this helicity frame based on the language of helicity amplitudes and Wigner-D function can be derived [BESIII:2019odb].
@@
The joint angular distribution of the single-side decay chains can be derived from polarization transmission [Chen:2019hqi] or equivalently the helicity density matrix [Perotti:2018wxm]. The angular distribution of the decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to\Lambda(\to p\pi^{-})\pi^{+})$ is shown below:
@@
where the $\alpha_{p\pi^{-}}$ denotes the decay asymmetry parameter in the weak hadronic decay $\Lambda\to p\pi^{-}$. The parameter $\Delta_{\Lambda\pi^{+}}$ is the phase shift between the helicity amplitudes of $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ satisfying $\beta_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\sin\Delta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\cos\Delta_{\Lambda\pi^{+}}$. Formulas of other two-body decays can be found in Ref. [BESIII:2019odb].
Although the analysis of the angular distribution of these four two-body CF channels provides effective constraints, the three-body channel $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, with a low background and large production rate, brings beneficial additional sensitivity. Benefiting from the transition amplitude analysis reported by the LHCb Collaboration [LHCb:2023crj], the angular distribution of decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to pK^{-}\pi^{+})$ can be expressed explicitly as:
@@
where $\phi_{1},\theta_{1},\phi_{2}$ are the helicity angles for the intermediate state ($K^{-}\pi^{+}$), $R_{ij}(\phi_{1},\theta_{1},\phi_{2})$ represents the 3-dimension $Z$-$Y$-$Z$ rotation matrix and $\kappa$ denotes a set of Dalitz-plot variables $(M(pK^{-})^{2},M(K^{-}\pi^{+})^{2})$. The $\vec{\alpha}(\kappa)$ and $I_{0}(\kappa)$ are model-agnostic representations for polarization dependence of the decay rate and the total differential decay rate over $\kappa$, respectively, which are taken from the $\Lambda_{c}^{+}$ polarimetry with the amplitude model both measured by the LHCb Collaboration [LHCb:2023crj, LHCb:2022sck].